OpenStudy (anonymous):

x2 – 81 = 0 solve the equation using square roots

OpenStudy (anonymous):

x^2 - 81 = 0 +81 +81 x^2=81 sqrt(x^2) = Sqrt(81) x=9

OpenStudy (anonymous):

hope I helped